OpenStudy (e28195):

For the function r(x) = 3/4cot(x) the x-intercepts occur at ____ radians. a. Πn b. Π/2 + Πn c. 3Π/4 + Πn d. Π/2

OpenStudy (e28195):

Please help

OpenStudy (jackthegreatest):

if you look at the graph, you can see that the x-intercept appears at every multiple of pi/2

OpenStudy (jackthegreatest):

so b

OpenStudy (jackthegreatest):

n is an integer, multiplying would create multiples of pi/2 (it could be negative)

OpenStudy (e28195):

Oh i see thank you!

OpenStudy (jackthegreatest):

np

OpenStudy (danjs):

\[\large r(x)=\frac{ 3 }{ 4 }*\cot(x) = \frac{ 3 }{ 4 }*\frac{ \cos(x) }{ \sin(x) }\] That will be zero when cos(x)=0. The top part of the fraction is zero. cos(x) = 0 for angles x= pi/2 and x= 3pi/2

OpenStudy (danjs):

or every pi radians starting with pi/2

OpenStudy (e28195):

oh that's a much more detailed response thanks :)